Exercise 14-12

On January 2, 2015, Stellar Corporation issued $1,550,000 of 10% bonds at 98 due December 31, 2024. Interest on the bonds is payable annually each December 31. The discount on the bonds is also being amortized on a straight-line basis over the 10 years. (Straight-line is not materially different in effect from the preferable “interest method.”)

The bonds are callable at 101 (i.e., at 101% of face amount), and on January 2, 2020, Stellar called $930,000 face amount of the bonds and redeemed them.

Ignoring income taxes, compute the amount of loss, if any, to be recognized by Stellar as a result of retiring the $930,000 of bonds in 2020. (Round answer to 0 decimal places, e.g. 38,548.)

Loss on redemption $enter a dollar amount of loss on redemption rounded to 0 decimal places


Prepare the journal entry to record the redemption. (Round answers to 0 decimal places, e.g. 38,548. If no entry is required, select "No Entry" for the account titles and enter 0 for the amounts. Credit account titles are automatically indented when amount is entered. Do not indent manually.)

Date

Account Titles and Explanation

Debit

Credit

January 2, 2020

enter an account title to record the transaction on January 2, 2017

enter a debit amount

enter a credit amount

enter an account title to record the transaction on January 2, 2017

enter a debit amount

enter a credit amount

enter an account title to record the transaction on January 2, 2017

enter a debit amount

enter a credit amount

enter an account title to record the transaction on January 2, 2017

enter a debit amount

enter a credit amount
Answer #1

Answer

--Loss on redemption = $ 18,600

A Face Value $930,000
B = A x 98/100 Issue Price $911,400
C = A - B Discount on Bonds Payable $18,600
D = 10 years No. of interest payment                                 10
E = C/D Annual amortisation of discount $1,860
F = E x 5 years Discount amortised till 2 Jan 2020 [5 years] $9,300
G = C - F Unamortised discount $9,300
H = A - G Book Value of Bonds Payable $920,700
I = A x 101% Cash paid on redepmtion $939,300
J = I - H Loss on redemption $18,600 Answer

--Journal entry to record redemption

Date Accounts title Debit Credit
02-Jan-20 Bonds Payable $930,000
Loss on redemption $18,600
   Discount on Bonds payable $9,300
   Cash $939,300